# A neural networks method to search for long transient gravitational waves

### v2

@@ IV.4 Denoiser improvements: masked loss
We need to increase the denoiser performance to reach better FAP or Eff. In particular, we need to improve our ability to retrieve the signal. During the training phase, we noticed that the loss decreased smoothly to a plateau. So in order to obtain better results, a change is needed, either in the NN structure or in the training procedure. A study of a diverse neural network pipeline will be the subject of future works. Here we report how we tried to highlight the structure of the signal and help the model learn it, using a masked loss. We note that the number of pixels in a single map (360 000) is three orders of magnitude higher than the number of pixels that form the signal($\sim$600). This suggests modifying the training scheme through the use of a mask in order to improve the learning process. Initially, each map was multiplied by a matrix that had 1 on the signal pixels and 0 otherwise. As the training proceeded, using a gaussian blurring function, we enlarged the area of the map that was not set to zero, increasing the difficulty of the task. A more detailed explanation of visual attention methods, such as masked loss, can be found in [62]. In Fig. (8), we report an example of the output of the denoiser. In particular, on the left-hand panel there is a map containing noise plus a signal; on the center panel there is the signal we want to retrieve; and finally, on the right hand panel there is the output of the denoiser.
@@
In Fig. (9), we report the ROC curves, obtained by the previous training technique. On the x-axis, there is the FAP, while on the y-axis, there is the Eff as a function of classifier threshold. The blue curve refers to learning without masked loss, i.e., what we have illustrated in Sec. (IV.3). The orange one is the best result that we have obtained after different trials with the masked loss, changing the parameters of the blurring. In Fig. (10b) we report the same quantities as in Fig. (6). We observe that the values are better:
$\mathrm{FAP}=2\%\ \ \ \ \mathrm{FAP}_{1800}=2\%$ (21)
$\mathrm{Eff}=90\%\ \ \ \ \mathrm{Eff}_{1800}=98\%.$
So the FAP decreases while the efficiency increases. At the same time, in Fig. (11) we report the distance of the source associated with the retrieved signals as a function of the parameter values. We see that even if the classifier identifies more signals, it still has problems in the upper right corner.
Figure 10: On the left panel, efficiency per signal (orange), efficiency per signal with a maximum frequency of 1800 Hz (red x), FAP (cyan), FAP with a maximum frequency of 1800 Hz (blue x) and F1 score (purple) are shown for different values of the threshold to discriminate, in each map, among the presence or absence of a signal. The straight fuchsia line corresponds to the maximum of the F1 score. On the right panel, a zoom in the range of thresholds between 0.4 and 0.7 is shown, with the FAP on the left y-axis, and the efficiency on the right y-axis
@@
$\mathrm{FAP}=4\%\ \ \ \ \mathrm{FAP}_{1800}=4\%$ (22)
$\mathrm{Eff}=92\%\ \ \ \ \mathrm{Eff}_{1800}=98\%.$
It is important to observe that we are conducting our analysis using data from a single interferometer. By combining data from a network of detectors, we expect to be able to lower the FAP. In fact, if we train our models separately on different interferometers data and then combine the results, we should lower our FAP down to the product of the different FAPs in the best-case scenario. This would allow us to choose a less stringent threshold to classify a map as having a signal.
@@ IV.5 Changing the braking index
To test the robustness of our method, we tried to test the denoiser and later train and test the classifier on the same signals, the same $f_{0}$ and $\epsilon$, but this time with a braking index randomly taken in the range [3.5, 5]. It is important to be able to retrieve signals with different braking indexes in order to take into account the different radiation processes of the star. In particular, we would like to have the capability of finding signals in regimes that differ from purely gravitational emissions. We are not training the denoiser again; we are using the best model of the previous subsection, trained with n=5. Then, with the new denoised maps, we trained the classifier again. The idea is to show that even if we make a restrictive assumption, the model is able to retrieve different power laws from the one we trained it with. In Fig. (12), we show the new parameter space, where $f_{0}$ and $\epsilon$ are the same as the previous section, while we changed $n$. For what concerns preserving the signal, in Fig. (13), we reported the best value of the overlap for each set of maps as a function of initial frequency, ellipticity and braking index.
@@
From the right-hand panel, we can notice that there is no trend in the braking index, in fact, the values of the overlap do not seem to depend on $n$. Although, on the left-hand panel, we notice the usual trend in $f_{0}$ and $\epsilon$. However, the different braking index makes the classification procedure easier in the upper right corner of the parameter space while worsening the situation in the middle, as we can notice comparing with Fig. (5).
As a consequence, the classification task is more difficult. In Fig. (14), we report the values of efficiency, FAP and F1 score for different thresholds; the maximum of the F1 score (fuchsia line) corresponds to
@@
The efficiency is worse than (20, 21), but we gain a better FAP, so fixing the FAP at $4\%$, as in the first case Eq. (20), we obtain:
@@ IV.6 Computational load
The total computing time for all the tasks described in this paper has been of the order of several hours. In particular, the creation of all the maps ($\sim$ 7000) on a M1 CPU took about 6 hours. In addition, on a single NVIDIA L40S GPU with 48GB of RAM, we needed 3 hours to train the denoiser with $\sim$2000 maps and 20 minutes to denoise $\sim$5000 maps. To train the classifier, 30 minutes were needed on $\sim$4000 maps and the classification task took less than 5 minutes on $\sim$1000 maps. In other words, to train both models, we needed $\sim$ 6000 maps of 1200 s that correspond up to 40 days of data. For more information, see the appendix (C).
@@ V Conclusions
The goal of this study was to develop machine learning techniques for the search of tGW signals, such as those emitted by newborn magnetars, in interferometric GW data.
We built, using a machine learning approach, a classifier to split time-frequency maps into two categories: the presence or the absence of signals. To help with the classification task, we built a denoiser.
To test the performance of our method, we simulated time-frequency maps containing noise plus signals. The noise was simualted according to the LIGO O4 design sensitivity curve, as explained in [48], and the signal according to the rotating ellipsoid model (e.g. [46]) with $n=5$.
Here we reported our results for signals having an initial amplitude of $2\times 10^{-23}$, the minimum value at which the use of the denoiser was beneficial to train the classifier.
We illustrate how this procedure can be used to search for tGWs emitted by newly born magnetars, and show that the denoiser plays a crucial role in the successful operation of the classifier. Indeed, we are not able to classify maps that have not been passed through the denoiser at an initial amplitude of $2\times 10^{-23}$. Moreover, the procedure proved to be robust to different power laws with $n\in[3.5,5]$, even if the denoiser was trained with n=5.
An extremely relevant point is the rapidity of this technique and the limited requirements in terms of computing power. In fact, the denoising procedure takes 3 hours for the training ($\sim$ 2000 maps) and 20 minutes for inference ($\sim$ 5000 maps). While the classification takes 30 minutes for the training ($\sim$ 4000 maps) and a few minutes to classify ($\sim$ 1000 maps). So with a trained pipeline, it is possible in about ten minutes to denoise and classify $\sim 1000$ maps. In other words, in a few minutes, we are able to analyze the equivalent of $\sim$14 days of data.
@@
In [30] a search for a tCW from a possible newly formed magnetar was conducted following the discovery of GW170817. In that case, different parameters were used for the simulation of the signal, in particular a moment of inertia that was three times bigger than the value we have used (at the limit of the mass ranges described in literature). As it is possible to notice in Eq. (7), the higher the momentum of inertia, the higher the amplitude, and the easier it is to classify signals.
A fair comparison of our method to the methods employed in [30] is not straightforward because the considered parameter space is different in the two cases, and there are also differences in the assumed NS moment of inertia. Overall, given the optimistic assumptions made in [30], we are confident that our technique will allow us to improve future searches to identify tCW candidates. More sensitive and model-dependent methods can be devised to perform a follow-up of signal candidates and measure signal parameters.
@@ Appendix C GPU usage and NN training
To speedup the NN training and inference times, we used an NVIDIA L40S GPU with 48GB of RAM, installed on the Virgo INFN Rome cluster. The denoiser for the first trail, Sec. (IV.2), was trained for fifty epochs, with a batch size of 8. Beyond that, the loss did not show significant improvement. After the implementation of the masked loss, Sec. (IV.4), we trained the model for 200 epochs. In the first 160, we used masked loss, changing the width of the blurring function every 16 epochs. In the last 40 epochs, we trained it using the whole map while computing the loss function. As in the previous case, we used as batch size 8.
For what it concerns the classifier, we trained for 40 epochs with a batch size of 40 in the first two scenarios, Secs. (IV.3,IV.4), and 60 epochs in the last one, Sec. (IV.5).

### v4

@@ IV.4 Denoiser improvements: masked loss
We need to increase the denoiser performance to reach better FAP or Eff. In particular, we need to improve our ability to retrieve the signal. During the training phase, we noticed that the loss decreased smoothly to a plateau. So in order to obtain better results, a change is needed, either in the NN structure or in the training procedure. A study of a diverse neural network pipeline will be the subject of future work. Here we report how we tried to highlight the structure of the signal and help the model learn it, using a masked loss. We note that the number of pixels in a single map (360 000) is three orders of magnitude higher than the number of pixels that form the signal ($\mathcal{O}$(600))  [^4] As we have already stated in Sec. III.2, we chose the size of the maps taking into consideration the properties of the signal we are interested in.. This suggests modifying the training scheme through the use of a mask in order to improve the learning process. Initially, each map was multiplied by a matrix that had 1 on the signal pixels and 0 otherwise. As the training proceeded, using a gaussian blurring function, we enlarged the area of the map that was not set to zero, increasing the difficulty of the task. A more detailed explanation of visual attention methods, such as masked loss, can be found in [62]. In Fig. 8, we report an example of the output of the denoiser. In particular, on the left-hand panel there is a map containing noise plus a signal; on the center panel there is the signal we want to retrieve; and finally, on the right hand panel there is the output of the denoiser.
@@
In Fig. 9, we report the ROC curves, obtained by the previous training technique. On the x-axis, there is the FAP, while on the y-axis, there is the Eff as a function of classifier threshold. The blue curve refers to learning without masked loss, i.e., what we have illustrated in Sec. IV.3. The orange one is the best result that we have obtained after different trials with the masked loss, changing the parameters of the blurring. In Fig. 10 we report the FAP and Eff in the masked loss case (brown and cornflower blue x), compared with the same quantities as in Fig. 6. We observe that the values are better:
$\mathrm{FAP}=2\%\ \ \ \ \mathrm{Eff}=90\%\$ (23)
So the FAP decreases while the efficiency increases. At the same time, in Fig. 11 we report the distance of the source associated with the retrieved signals as a function of the parameter values. We see that even if the classifier identifies more signals, it still has problems in the upper right corner.
[…]
@@
$\mathrm{FAP}=4\%\ \ \ \ \mathrm{Eff}=92\%$ (24)
as shown in Fig. 10, in correspondence of the green vertical line. It is important to observe that we are conducting our analysis using data from a single interferometer. By combining data from a network of detectors, we expect to be able to lower the FAP. In fact, if we train our models separately on different interferometers data and then combine the results, we should lower our FAP down to the product of the different FAPs in the best-case scenario. This would allow us to choose a less stringent threshold to classify a map as having a signal.
@@ IV.5 Changing the braking index
At the beginning of our work we trained our model assuming a braking index $n=5$ (see equation (1)), that is, without taking into account the electromagnetic contribution to the spin-down (which is described, for instance, in [63]). To assess the robustness of the model, we tested its performance by injecting signals with different values of $n$, since our goal in the future is to generalize our technique by combining electromagnetic and gravitational torques.
So using the models already trained, we tried to test the denoiser and later train and test the classifier on signals, with the same $f_{0}$ and $\epsilon$, with n randomly taken in the range [3.5, 5]. It is important to be able to retrieve signals with different braking indexes in order to take into account the different emission processes of the star. In particular, we would like to have the capability of finding signals in regimes that differ from purely gravitational emissions. We did not train the denoiser again; rather we have used the best model of the previous subsection, trained with $n$ = 5. Then, with the new denoised maps, we trained the classifier again. The idea is to show that even if we make a restrictive assumption, the model is able to retrieve different power laws from the one we trained it with. In Fig. 12, we show the new parameter space, where $f_{0}$ and $\epsilon$ are the same as the previous section, while we changed $n$. For what it concerns the fraction of the pixels of the signal which is preserved, in Fig. 13, we report the best value of the overlap for each set of maps as a function of initial frequency, ellipticity and braking index.
@@
From the right-hand panel, we can notice that there is no trend in the braking index, in fact, the values of the overlap do not seem to depend on $n$. Nevertheless, on the left-hand panel, we notice the usual trend in $f_{0}$ and $\epsilon$. However, the different braking index makes the classification procedure easier in the upper right corner of the parameter space while worsening the situation in the middle, as we can notice comparing with Fig. 5.
As a consequence, the classification task is more difficult. In Fig. 14, we report the values of efficiency, FAP and F1 score for different thresholds; the maximum of the F1 score (fuchsia line) corresponds to
@@
Figure 14: Same quantities as in Fig. 6. The denoiser has been trained with the masked loss on signal with $n$ = 5. The classified signals have n $\in[3.5,5]$.
The efficiency is worse than (22, 23), but we gain a better FAP, so fixing the FAP at $4\%$, as in the first case Eq. (22), we obtain:
@@ IV.6 Computational load
The total computing time for all the tasks described in this paper has been of the order of several hours. In particular, the creation of all the maps ($\mathcal{O}$(7000)) took about 6 hours. In addition, we needed 3 hours to train the denoiser with $\mathcal{O}$(2000) maps and 20 minutes to denoise $\mathcal{O}$(5000) maps. To train the classifier, 30 minutes were needed on $\mathcal{O}$(4000) maps, and the classification task took less than 5 minutes on $\mathcal{O}$(1000) maps. In other words, to train both models, we needed $\mathcal{O}$(6000) maps of 1200 s that correspond up to 40 days of data. For more information, see Appendix C.
IV.7 Discussion
In [30] a search for a tCW from a possible newly formed magnetar was conducted following the discovery of GW170817. In that case, different parameters were used for the simulation of the signal, in particular a moment of inertia that was three times bigger than the value we have used (at the limit of the mass ranges described in literature). As it is possible to notice in Eq. (7), the higher the moment of inertia, the higher the amplitude, and the easier it is to classify signals.
A fair comparison of our method to the methods employed in [30] is not straightforward because the considered parameter space is different in the two cases, and there are also differences in the assumed NS moment of inertia. However, we ran a code to estimate the optimal sensitivity of the generalized Frequency Hough (GFH) pipeline (see [25], Eq. (34)). We used the same signals, under the same simulated $O4$ noise. We obtain a maximum distance that is slightly less than twice that obtained by our method. It should be kept in mind, that this result is obtained under ideal conditions, in which the entire signal is fully integrated into the GFH calculation and the optimal data segment duration is used in all the parameter space. This approach is not realistically applicable for a general analysis in which signal parameters are not known a priori. A systematic comparison of these two methods sensitivities is beyond the scope of this article. Overall, we are confident that our technique will allow us to improve future searches to identify tCW candidates.
@@ V Conclusions
The goal of this study was to develop machine learning techniques for the search of tCW signals, such as those emitted by newborn magnetars, in interferometric data. We built, using a machine learning approach, a classifier to split time-frequency maps into two categories: presence or absence of signals. To help with the classification task, we built a denoiser.
To test the performance of our method, we simulated time-frequency maps containing noise plus signals. The noise was simulated according to the LIGO O4 design sensitivity curve, as explained in [48], and the signal according to the rotating ellipsoid model (e.g. [45]) with $n=5$.
Here we report our results for signals having an initial amplitude of $2\times 10^{-23}$, the minimum value at which the use of the denoiser was beneficial to train the classifier.
We illustrate how this procedure can be used to search for tCWs emitted by newly born magnetars, and show that the denoiser plays a crucial role in the successful operation of the classifier. Indeed, we are not able to classify maps that have not been passed through the denoiser at an initial amplitude of $2\times 10^{-23}$. Moreover, the procedure proved to be robust to different power laws with $n\in[3.5,5]$, even if the denoiser was trained with n=5.
An extremely relevant point is the rapidity of this technique and the limited requirements in terms of computing power. In fact, we are able to analyze the equivalent of $\mathcal{O}$(14) days of data in a few minutes.
@@
Overall, the proposed method has a sensitivity which is comparable with the already existing methods. In general, more sensitive and model-dependent methods can be devised to perform a follow-up of signal candidates and measure signal parameters.
@@ Appendix C GPU usage and NN training
To speedup the NN training and inference times, we used an NVIDIA L40S GPU with 48GB of RAM, installed on the Virgo INFN Rome cluster. The denoiser for the first train, Sec. IV.2, was trained for fifty epochs, with a batch size of 8. Beyond that, the loss did not show significant improvement. After the implementation of the masked loss, Sec. IV.4, we trained the model for 200 epochs. In the first 160, we used masked loss, changing the width of the blurring function every 16 epochs. In the last 40 epochs, we trained it using the whole map while computing the loss function. As in the previous case, we used as batch size 8.
Concerning the classifier, we trained for 40 epochs with a batch size of 40 in the first two scenarios, Secs. IV.3,IV.4, and 60 epochs in the last one,Sec. IV.5.